OpenStudy (anonymous):

when does the sign switch for an equality problem? when does >= switch to <=?

zepdrix (zepdrix):

When you `multiply` or `divide` a value across the inequality.

zepdrix (zepdrix):

Example:\[\Large\rm -3x\le 6\]Dividing by -3 means we have to also flip the sign,\[\Large\rm x\ge \frac{6}{-3}\] \[\Large\rm x\ge-2\]

OpenStudy (anonymous):

so for this one would it stay the same -2(x+3) >= 8?

zepdrix (zepdrix):

In order to solve for x, a good first step would be to divide both sides by -2.

zepdrix (zepdrix):

In that case, we're dividing a negative across an inequality, yes?

zepdrix (zepdrix):

So noooo :O it wouldn't stay the same. It would flip, yah? :)

OpenStudy (anonymous):

yaaassss thank you!

OpenStudy (anonymous):

might have more questions for you!